OpenStudy (anonymous):

What are the x- and y- intercepts of the following equation: y = -0.05x + 16 ????

OpenStudy (aravindg):

\[\huge \text{WELCOME TO OPENSTUDY !}\]

OpenStudy (aravindg):

Try to write the equation in form x/a+y+b=1

OpenStudy (aravindg):

then a will be x intercept and b will be y intercept

OpenStudy (anonymous):

x intercept is the point that lies on the x-axis and is when y=0 y intercept is the point that lies on the y-axis and is when x=0

OpenStudy (anonymous):

thank you :]

OpenStudy (anonymous):

the y intercept is what you get if \(x=0\) your equation is written in the "slope intercept" form \(y=mx+b\)so the y intercept is displayed for you as \(b\) in your cases \(16\)

OpenStudy (anonymous):

so it would be (0, 16), (32, 0) ?

OpenStudy (anonymous):

the first one is right, i don't think the second one is if it is \(.05\) and not \(.5\) then you are off by a factor of 10

OpenStudy (anonymous):

\[ y = -0.05x + 16 \]\[0=-.05x+16\] \[.05x=16\] \[x=16\div .05=1600\div 5=320\]

OpenStudy (anonymous):

(320, 0) ?

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

Solve 3(x - 2) > 4(2x + 11). can u help me with this one please ?